OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4 Someone please show me how to do this?

OpenStudy (smileyxl3):

Expand it. 3 x+3 = -2 x-2 After that, you would get x = -1

OpenStudy (anonymous):

Ok thank you :)

OpenStudy (smileyxl3):

You're welcome. You can click the Best Response if the answer helped :)

OpenStudy (anonymous):

Would you be able to possibly help me with something else? Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers. I have no idea how to do this

OpenStudy (smileyxl3):

Two-step equation? Is it for the previous equation to get to the answer?

OpenStudy (anonymous):

No its a completely different problem

OpenStudy (smileyxl3):

There isn't a equation to go with it? Or you supposed to make up your own equation?

OpenStudy (anonymous):

You are supposed to make your own equation and show how you got to it, but I have no idea what a two step equation with all real numbers is supposed to be.

OpenStudy (smileyxl3):

Ok. Example equation 4x + 2 = 4x + 2 → 4x = 4x → x = x

OpenStudy (anonymous):

Ok....so is that an example of a two step equation? (i hate algebra)

OpenStudy (anonymous):

Thank you

OpenStudy (smileyxl3):

Pretty sure that is a two step equation. If you wanted to add a bit more, you could add that you subtract 2 from each side but that's about it.

OpenStudy (anonymous):

Thank you ^_^